A cash flow at time zero (now) of $14,551 is equivalent to another cash flow that is an EOY annuity of $2,700 over seven years (starting at year 1). Each of these two cash-flow series is equlvaient to a thied sene which is a uniform gradient series What is the value of G for this third series over the same seven-year ime interval? Assume that the cash flow at the end of year one is zero. Choose the comect answer below O A. $989 O B. $77 O C. $296 O D. $1,450 O E. Not enough information given
Answer #1

14551 = 2700*(P/A, i, 7)

14551 / 2700 = (P/A, i, 7)

5.389 = (P/A, i , 7)

(** from compound interest table we find that (P/A) at n = 7 is 5.389 when i = 7%

so,

i= 7%

14551 = G*(P/G, 7%, 7)

14551 = G* 14.715

G = 988.85 = 989

Ans is a) 989.